If the mean of a normal distribution is 70 and the standard deviation is 10 what is the probability the item selected at random for this sample will be greater than 80?
Answer by stanbon(75887)

If the mean of a normal distribution is 70 and the standard deviation is 10 what is the probability the item selected at random for this sample will be greater than 80?
------
z(80) = (80-70)/10 = 1
----
P(x > 80) = P(z > 1) = 0.1587
===================
Cheers,
Stan H.